Boundary Value Problem for Functional Differential Equation

Mustapha Yebdri University of Tlemcen, Algeria

A suitable periodic boundary conditions for a functional differential equations ẋ(t) = f(t, x, x_{t}) are conditions of the form x_{0}(θ) = x_{2π}(θ). In this talk we use the notion of upper and lower solutions coupled with monotone iterative method to prove the existence of solutions of this periodic boundary value problem.